OpenStudy (anonymous):

Simplify: (7−3)÷2+4*6

OpenStudy (anonymous):

@LynFran

OpenStudy (arindameducationusc):

Use BODMAS

OpenStudy (arindameducationusc):

7-3=4

OpenStudy (arindameducationusc):

4/2=2

OpenStudy (arindameducationusc):

and 6*4=24

OpenStudy (arindameducationusc):

now 2+24, you get the answer, Got it?

OpenStudy (naga447):

=-22

OpenStudy (arindameducationusc):

???? @naga447 You okay?

OpenStudy (naga447):

7-2-2-4*6= 4-2-4*6= 2-4*6= 2-24= =-22

OpenStudy (naga447):

If you take 24 from 2 you get a negative

OpenStudy (arindameducationusc):

@naga447 Thats not right....

OpenStudy (arindameducationusc):

@jcoury could you please check this......

OpenStudy (anonymous):

that looks fine